OpenStudy (fibonaccichick666):

Just need a quick check on a geometry proof: Let O be the inscribed circle in the isosceles trapezoid ABCD where AB is parallel to CD, AB=a, and CD=b. Find the radius of the inscribed circle. Pf: \(r=\frac{h}{2}=\frac{\sqrt {ab}} {2}\) In order for there to be an inscribed circle, we know that \(AD=BD=\frac{a+b}{2}\). Then we can get h by the pythagorian thm as \(h=\sqrt{ab}\). Then because the inscribed circle must be tangent to both the top and the bottom of the trapezoid, we know that a diameter is parallel to the height, so our radius is h/2. What do you think?

OpenStudy (fibonaccichick666):

Here it is since the math won't show: \(r=\frac{h}{2}=\frac{\sqrt {ab}} {2}\) In order for there to be an inscribed circle, we know that \(AD=BD=\frac{a+b}{2}\). Then we can get h by the pythagorian thm as \(h=\sqrt{ab}\). Then because the inscribed circle must be tangent to both the top and the bottom of the trapezoid, we know that a diameter is parallel to the height, so our radius is h/2.

OpenStudy (fibonaccichick666):

My picture as well: 

OpenStudy (fibonaccichick666):

I did not mark the angles because I didn't use that they are congruent

OpenStudy (fibonaccichick666):

@ganeshie8 , do you mind taking a quick look?

OpenStudy (fibonaccichick666):

or could you take a look @mathmale ?

OpenStudy (fibonaccichick666):

More than anything, I just need to know if I need to expand on any points.

OpenStudy (mathmale):

Your approach looks quite thorough, and you're to be congratulated for that. Regarding "we know that a diameter is parallel to the height, so our radius is h/2": I agree. Would you mind demonstrating how you'd use the Pyth. Thm. to support your statement, "In order for there to be an inscribed circle, we know that AD=BD=a+b2. Then we can get h by the pythagorian thm as h=ab−−√. "

ganeshie8 (ganeshie8):

\(AD=BD=\frac{a+b}{2} \)

ganeshie8 (ganeshie8):

thats a typo right ?

ganeshie8 (ganeshie8):

it should be : \(AD=BC=\frac{a+b}{2}\)

ganeshie8 (ganeshie8):

overall it looks good to me !

OpenStudy (fibonaccichick666):

Thank you, I just didn't show pythag here, but I did list it on the paper I'm using: \[h^2=AD^2-DP^2\]Forgot to name P on my pic, but AP is the perpendicular. \[h^2=[\frac{a+b}{2}]^2- [\frac{b-a}{2}]^2\] \[=\frac{a^2+2ab+b^2-b^2+2ab-a^2}{4}\] \[=\frac{4ab}{4}\] \[h^2=ab\] \[h=\sqrt{ab}\] @ganeshie8 it just looks like a minus, it is a plus

OpenStudy (fibonaccichick666):

ohhh yea type o

OpenStudy (fibonaccichick666):

It should be BC

ganeshie8 (ganeshie8):

nice work !

OpenStudy (fibonaccichick666):

Thank you, I just needed to check it. I am absolutely awful at knowing if a proof is detailed enough

OpenStudy (fibonaccichick666):

Thanks to both of you, I really appreciate the help!

ganeshie8 (ganeshie8):

yes, it is not detailed enough thats the reason u got many questions.... guess u need to justify few things in the proof

ganeshie8 (ganeshie8):

unless they're not proven/used everyday in ur life

OpenStudy (fibonaccichick666):

yea, I actually have both of the questions written on my paper as side work, I'll just include those as well

ganeshie8 (ganeshie8):

for example, nobody remembers on top of their heads : AD = BC = (a+b)/2

OpenStudy (fibonaccichick666):

ok, so I'll elaborate on why there

ganeshie8 (ganeshie8):

okay !

OpenStudy (fibonaccichick666):

Thanks again!

ganeshie8 (ganeshie8):

np :)

OpenStudy (phi):

I think I would draw it like this